9.A company’s dividend grows at a constant rate of 4 percent p.a.. Last week it paid a dividend of $6.95. If the required rate of return is 18 percent p.a., what is the price of the share 4 years from now? (round to nearest cent)

Select one:

a. $60.40

b. $31.15

c. $58.08

d. $100.10

10.Which of the following best describes the constant-growth dividend discount model?

Select one:

A. It is the formula for the present value of a growing annuity

B. It is the formula for the present value of an ordinary annuity.

C. It is the formula for the present value of a finite, uneven cash flow stream

D. It is the formula for the present value of a growing perpetuity
Answer #1

9. Calculation of price of the share 4 years from now:

Given Growth rate(g) = 4%, Current Dividend(d0) = $6.95, Ke= 18%

Price after 4 years(P4) = Dividend in the 5th year(d5) / Ke-g

= $8.4557/(0.18-0.04)

= $60.40 (rounded off)

W.N 1: Calculation of dividend in the 5th year

d5  = d0(1+g)5 = $6.95*(1+0.04)5 = $8.4557

so correct option is (a).

10. Constant growth dividend discount model assumes company exists forever and pays dividend per share that increase at a constant rate. It is also called Gordon growth model. It is the formula for the present value of a growing perpetuity.

P0 = d1/ke-g

where p0 = price of stock

d1 = estimated dividend for next period

ke = required rate of return

g = growth rate

so correct option is (d).